# A simple lower bound for the complexity of estimating partition functions on a quantum computer

### v1

@@ Abstract
We study the complexity of estimating the partition function $\mathsf{Z}(\beta)=\sum_{x\in\chi}e^{-\beta H(x)}$ for a Gibbs distribution characterized by the Hamiltonian $H(x)$. We provide a simple and natural lower bound for quantum algorithms that solve this task by relying on reflections through the coherent encoding of Gibbs states. Our primary contribution is a $\Omega(1/\epsilon)$ lower bound for the number of reflections needed to estimate the partition function with a quantum algorithm. We also prove a $\Omega(1/\epsilon^{2})$ query lower bound for classical algorithms. The proofs are based on a reduction from the problem of estimating the Hamming weight of an unknown binary string.
@@ 1 Introduction
Besides the applications in physics or statistical mechanics [GZ04, TMCA22, AKKR02], Gibbs distributions and the partition functions are also used in mathematics and computer science. Examples of applications of partition functions include the convergence analysis of stochastic gradient descent [SSJ20, Shi21], counting proper $k$-colorings of a graph [Jer95, Vig99, BC20], the ferromagnetic Ising model [MS13, WTR+20], as well as counting matchings and independent sets [HW20, AHN+22, CH23]. Some of these applications depend on efficiently estimating the value of the partition function at different temperatures. While it is relatively easy to estimate at high temperatures, estimation becomes more challenging at low temperatures, although the low-temperature regime is where the aforementioned applications are found. For instance, the number of proper $k$-colorings of a graph [Jer95, ŠVV09] and the volume of a convex body [DF91, CCH+23] are obtained from the partition function at very low temperatures.
@@
The literature contains many algorithms, both classical and quantum, to estimate partition functions. The majority of these approaches employ a strategy that begins in a high-temperature region and progressively lowers the temperature, using a telescoping product of partition functions at neighboring temperatures. Initial studies primarily employed a non-adaptive cooling schedule, implying that the gaps between neighboring temperatures are constant [DF91, BŠVV08]; subsequently, adaptive schedules were proposed [ŠVV09, Hub15, Mon15, Kol18, HW20, AHN+22, CH23].
@@
Most studies focus on improving the upper bound of partition function estimation. The study of lower bounds on the complexity of estimating partition function has not received as much attention. Exceptions are [ŠVV09, Kol18, CLL+22], which, however, do not directly address the question of providing a quantum lower bound for the problem in a general fashion: we discuss the lower bounds of these papers in a subsequent section, after introducing the necessary notation. The derivation of a general lower bound was also stated as an open question in [AHN+22]. Although a quantum lower bound of $\Omega(1/\epsilon)$ and a classical lower bound of $\Omega(1/\epsilon^{2})$ seem natural and almost unavoidable in this setting, to the best of our knowledge a formal proof is not discussed in the literature. It is possible that such results are part of the folklore: in that case, we welcome readers who are aware of existing proofs to contact us so that we can give proper attribution.
@@
In this note, using a simple argument, we establish a $\Omega(1/\epsilon)$ lower bound on the number of reflections needed by a quantum algorithm for estimating partition functions at low temperature. Existing quantum algorithms [Mon15, HW20, AHN+22, CH23] match this scaling in $\epsilon$. With a similar argument, we also establish a $\Omega(1/\epsilon^{2})$ query lower bound for classical algorithms for estimating partition functions; the fastest classical algorithms [ŠVV09, Kol18] match this scaling in $\epsilon$.
@@ 1.2 Basic definitions
Let $\chi$ be a finite set and $H:\chi\to\{0,1,\cdots,n\}$ a function called the Hamiltonian. (The ground set for the Hamiltonian is often called $\Omega$ in the literature, but we use $\chi$ because we reserve $\Omega$ for the asymptotic notation for lower bounds.) The Gibbs distribution is defined as
@@
A measurement in the computational basis from (2) yields a sample from the Gibbs distribution (1). These qsamples are used in existing quantum algorithm to estimate $\mathsf{Z}(\beta)$, see, e.g., [Mon15, HW20, AHN+22, CH23].
@@ 1.3 Existing lower bounds
For classical algorithms, the first lower bound on partition function estimation is due to [ŠVV09]. [ŠVV09] gives a lower bound on the number of temperatures of a non-adaptive cooling schedule, and therefore this lower bound only applies to the class of simulated annealing algorithms with non-adaptive cooling schedules. [Kol18] establishes a lower bound of $\varOmega\left(\frac{\log{|\chi|}}{\epsilon^{2}}\right)$ for the estimation of partition function, in a computational model where we receive samples from $H(x)$ but we are now allowed to know the corresponding $x$, or query $H(x)$ for specific $x$. Our lower bound is not as strong, but it uses the more traditional query model and it has simple proof.
Regarding quantum algorithms, the only lower bound that we are aware of is [CLL+22], which gives a lower bound of $\varOmega(\epsilon^{-\frac{1}{1+4/k}})$ for the quantum query complexity of estimating $\int_{\mathbb{R}^{k}}e^{-f(x)}\textrm{d}x$, under the assumption that $f$ is $1.5$-smooth and $0.5$-strongly convex. This lower bound is different from ours in that we consider a general Hamiltonian defined over a discrete set, whereas [CLL+22] considers $f$ defined over real space and strongly convex. Note that a classical lower bound of $\varOmega(k/\epsilon^{2})$ for the same setting as [CLL+22] is proven in [GLL20].
@@ 1.4 Contributions
Our main contribution (Theorem 1) is a quantum query complexity lower bound for estimating partition functions with a simple proof. To state the lower bound we need to identify a meaningful “unit of work” for quantum algorithms to estimate $\mathsf{Z}(+\infty)$. The basic steps that [HW20, AHN+22, CH23] have in common is the reflection through the Gibbs states (2) at different inverse temperatures $\beta$, therefore our lower bound is stated in terms of the number of reflection through these states. Note that the quantum algorithms described in the above papers access the Hamiltonian only through these reflections.
@@ 2.1 Preliminaries
Proposition 1 (Quantum query tight bound for the Hamming weight problem [NW99] [CLL+22]).
@@
Let $m\in\left\{\ell,\ell^{\prime}\right\}$ be such that $\left|\frac{N}{2}-m\right|$ is maximized, and let $\Delta=\ell^{\prime}-\ell$. Then, given the quantum query oracle
@@ 2.2 Construction of the Hamiltonian
The restriction that reflections are invoked at most ${\cal O}(\log^{2}|\chi|/\epsilon)$ times is technical: known quantum algorithms for partition functions, e.g., [Mon15, HW20, AHN+22, CH23], achieve this scaling (possibly up to polylogarithmic factors, which we neglect for simplicity) or better.
@@
We first give an upper bound for the number of queries to $O_{H}$ to construct $\lvert\mu_{+\infty}\rangle$ using Proposition 2, and subsequently, we prove that with the same cost we can construct $\lvert\mu_{\beta}\rangle$ for any other $\beta$. For now we assume that $\|b\|_{1}$ is known, and later we show that an unknown value of $\|b\|_{1}$ has negligible impact. Our initial state $\lvert s\rangle$ is the uniform superposition of all states $\lvert x\rangle$:
@@
One potential issue still remains regarding the construction of a reflection via state preparation: we do not know $\|b\|_{1}$, therefore the state preparation above cannot be exact. We show that the overlap between the ideal state (assuming knowledge of $\|b\|_{1}$) and the imperfect state (assuming no knowledge of $\|b\|_{1}$) is ${\cal O}(\delta)$, and this does not significantly affect the final result.
@@ 4 Conclusions
In this paper we prove a quantum lower bound on the number of reflections for estimating partition function at low temperatures. This implies the optimality in the $\epsilon$ parameter of recent quantum algorithms for this problem [Mon15, HW20, AHN+22, CH23]. Our classical lower bound is weaker than the one in [Kol18], but it uses a more traditional computational model and has a simple proof.
The unresolved question is whether we can achieve a lower bound stronger than $\varOmega({1}/{\epsilon})$ for quantum algorithms, in particular incorporating $|\chi|$ into the lower bound. We attempted several constructions for $H$ and $\chi$, and the one discussed here is the most robust among those that we examined. It is possible that a different approach – one that does not uses a reduction from the problem of determining the Hamming weight of an unknown binary string – is necessary; in particular, a reduction from a problem with multiple instance parameters involved in the lower bound might be more successful in incorporating $|\chi|$.

### v2

@@ Abstract
We study the complexity of estimating the partition function $\mathsf{Z}(\beta)=\sum_{x\in\chi}e^{-\beta H(x)}$ for a Gibbs distribution characterized by the Hamiltonian $H(x)$. We provide a simple and natural lower bound for quantum algorithms that solve this task by relying on reflections through the coherent encoding of Gibbs states. Our primary contribution is a $\varOmega(1/\epsilon)$ lower bound for the number of reflections needed to estimate the partition function with a quantum algorithm. The proof is based on a reduction from the problem of estimating the Hamming weight of an unknown binary string.
@@ 1 Introduction
Besides the applications in physics or statistical mechanics [GZ04, TMCA22, AKKR02], Gibbs distributions and the partition functions are also used in mathematics and computer science. Examples of applications of partition functions include the convergence analysis of stochastic gradient descent [SSJ20, Shi21], counting proper $k$-colorings of a graph [Jer95, Vig99, BC20], the ferromagnetic Ising model [MS13, ${}^{+}$20], as well as counting matchings and independent sets [HW20, ${}^{+}$22, CH23]. Some of these applications depend on efficiently estimating the value of the partition function at different temperatures. While it is relatively easy to estimate at high temperatures, estimation becomes more challenging at low temperatures, although the low-temperature regime is where the aforementioned applications are found. For instance, the number of proper $k$-colorings of a graph [Jer95, ŠVV09] and the volume of a convex body [DF91, ${}^{+}$23] are obtained from the partition function at very low temperatures.
@@
The literature contains many algorithms, both classical and quantum, to estimate partition functions. The majority of these approaches employ a strategy that begins in a high-temperature region and progressively lowers the temperature, using a telescoping product of partition functions at neighboring temperatures. Initial studies primarily employed a non-adaptive cooling schedule, implying that the gaps between neighboring temperatures are constant [DF91, BŠVV08]; subsequently, adaptive schedules were proposed [ŠVV09, Hub15, Mon15, Kol18, HK20, HW20, ${}^{+}$22, CH23].
@@
Most studies focus on improving the upper bound of partition function estimation. The study of lower bounds on the complexity of estimating partition function has not received as much attention. Exceptions are [ŠVV09, Kol18, ${}^{+}$22, HK20], which, however, do not directly address the question of providing a quantum lower bound for the problem in a general fashion: we discuss the lower bounds of these papers in a subsequent section, after introducing the necessary notation. The derivation of a general lower bound was also stated as an open question in [${}^{+}$22]. Although a quantum lower bound of $\varOmega(1/\epsilon)$ seems natural and almost unavoidable in this setting, to the best of our knowledge a formal proof is not discussed in the literature. It is possible that such a result is part of the folklore: in that case, we welcome readers who are aware of existing proofs to contact us so that we can give proper attribution.
@@
In this note, using a simple argument, we establish a $\varOmega(1/\epsilon)$ lower bound on the number of reflections needed by a quantum algorithm for estimating partition functions at low temperature. Existing quantum algorithms [Mon15, HW20, ${}^{+}$22, CH23] match this scaling in $\epsilon$. With a similar argument, we also establish a $\varOmega(1/\epsilon^{2})$ query lower bound for classical algorithms for estimating partition functions; this is weaker than the lower bound of $\varOmega(\log|\chi|/\epsilon^{2})$ samples from the Hamiltonian recently given in [HK20], but we mention it because the proof is considerably simpler. The fastest classical algorithms [Kol18, HK20] match this scaling in $\epsilon$.
@@ 1.2 Basic definitions
Let $\chi$ be a finite set and $H:\chi\to\{0,1,\cdots,n\}$ a function called the Hamiltonian. (The ground set for the Hamiltonian is often called $\varOmega$ in the literature, but we use $\chi$ because we reserve $\varOmega$ for the asymptotic notation for lower bounds.) The Gibbs distribution is defined as
@@
A measurement in the computational basis from (2) yields a sample from the Gibbs distribution (1). These qsamples are used in existing quantum algorithm to estimate $\mathsf{Z}(\beta)$, see, e.g., [Mon15, HW20, ${}^{+}$22, CH23].
@@ 1.3 Existing lower bounds
For classical algorithms, the first lower bound on partition function estimation is due to [ŠVV09]. [ŠVV09] gives a lower bound on the number of temperatures of a non-adaptive cooling schedule, and therefore this lower bound only applies to the class of simulated annealing algorithms with non-adaptive cooling schedules. [Kol18] establishes a lower bound of $\varOmega\left(\frac{\log{|\chi|}}{\epsilon^{2}}\right)$ for the estimation of partition function, in a computational model where we receive samples from $H(x)$ but we are now allowed to know the corresponding $x$, or query $H(x)$ for specific $x$. [HK20] extends this to a lower bound in the black-box sampling model, showing optimality of existing classical algorithms. (Our classical lower bound is not as strong, but it has a very simple proof.)
Regarding quantum algorithms, the only lower bound that we are aware of is [${}^{+}$22], which gives a lower bound of $\varOmega(\epsilon^{-\frac{1}{1+4/k}})$ for the quantum query complexity of estimating $\int_{\mathbb{R}^{k}}e^{-f(x)}\textrm{d}x$, under the assumption that $f$ is $1.5$-smooth and $0.5$-strongly convex. This lower bound is different from ours in that we consider a general Hamiltonian defined over a discrete set, whereas [${}^{+}$22] considers $f$ defined over real space and strongly convex. Note that a classical lower bound of $\varOmega(k/\epsilon^{2})$ for the same setting as [${}^{+}$22] is proven in [GLL20].
@@ 1.4 Contributions
Our main contribution (Theorem 1) is a quantum query complexity lower bound for estimating partition functions with a simple proof. To state the lower bound we need to identify a meaningful “unit of work” for quantum algorithms to estimate $\mathsf{Z}(+\infty)$. The basic steps that [HW20, ${}^{+}$22, CH23] have in common is the reflection through the Gibbs states (2) at different inverse temperatures $\beta$, therefore our lower bound is stated in terms of the number of reflection through these states. Note that the quantum algorithms described in the above papers access the Hamiltonian only through these reflections.
@@
The classical lower bound is weaker than the lower bound of $\varOmega({\log|\chi|/\epsilon^{2}})$ given in [HK20], but the proof is simple. The results of [HK20] are very recent and we were not aware of them when working on our result, hence we did not try to quantize them. It is possible that their construction yields a stronger quantum lower bound than Thm. 1.
@@ 2.1 Preliminaries
Proposition 1 (Quantum query tight bound for the Hamming weight problem [NW99] [${}^{+}$22]).
@@
Let $m\in\left\{\ell,\ell^{\prime}\right\}$ be such that $\left|\frac{N}{2}-m\right|$ is maximized, and let Δ=normal-$\Delta=\ell^{\prime}-\ell$. Then, given the quantum query oracle
@@ 2.2 Construction of the Hamiltonian
The restriction that reflections are invoked at most ${\cal O}(\log^{2}|\chi|/\epsilon)$ times is technical: known quantum algorithms for partition functions, e.g., [Mon15, HW20, ${}^{+}$22, CH23], achieve this scaling (possibly up to polylogarithmic factors, which we neglect for simplicity) or better.
@@
We first give an upper bound for the number of queries to $O_{H}$ to construct $\lvert\mu_{+\infty}\rangle$ using Proposition 2, and subsequently, we prove that with the same cost we can construct $\lvert\mu_{\beta}\rangle$ for any other $\beta$. For now we assume that $\|b\|_{1}$ is know, and later we show that an unknown value of $\|b\|_{1}$ has negligible impact. Our initial state $\lvert s\rangle$ is the uniform superposition of all states $\lvert x\rangle$:
@@
One potential issue still remains regarding the construction of a reflection via state preparation: we do not know $\|b\|_{1}$, therefore the state preparation above cannot be exact. We show that the overlap between the ideal state (assuming knowledge of $\|b\|_{1}$) and the imperfect state (assuming no knowledge of $\|b\|_{1}$) is $1-{\cal O}(\delta)$, and this does not significantly affect the final result.
@@ 3.2 Proof
As previously remarked, [HK20] gives a stronger sampling lower bound than the one we provide above.
@@ 4 Conclusions
In this paper we prove a quantum lower bound on the number of reflections for estimating partition function at low temperatures. This implies the optimality in the $\epsilon$ parameter of recent quantum algorithms for this problem [Mon15, HW20, ${}^{+}$22, CH23]. Our classical lower bound is weaker than the one in [Kol18], but it uses a more traditional computational model and has a simple proof; [HK20] supersedes [Kol18] and shows a stronger lower bound in the standard black-box sampling computational model.
The unresolved question is whether we can achieve a lower bound stronger than $\varOmega({1}/{\epsilon})$ for quantum algorithms, in particular incorporating $|\chi|$ into the lower bound. We attempted several constructions for $H$ and $\chi$, and the one discussed here is the most robust among those that we examined. It is possible that a different approach – one that does not uses a reduction from the problem of determining the Hamming weight of an unknown binary string – is necessary; in particular, a reduction from a problem with multiple instance parameters involved in the lower bound might be more successful in incorporating $|\chi|$. It is possible that the approach of [HK20] leads to a stronger quantum lower bound: we did not know their result (which extends [Kol18]) until posting a first version of this manuscript on arXiv, hence we did not make any attempt at quantizing it yet.
Acknowledgements
We are grateful to D. Harris for pointing out [HK20] to us. G. Nannicini is partially supported by ONR award # N000142312585.